Cheryl Montoya picked up the phone and called her boss, Wes Chan, the vice president of marketing at Piedmont Fasteners Corporation: Wes, Im not sure how to go about answering the questions that came up at the meeting with the president yesterday. Whats the problem? The president wanted to know the break-even point for each of the companys products, but I am having trouble figuring them out. Tm sure you can handle it, Cheryl. And, by the way, I need your analysis on my desk tomorrow morning at 8:00 sharp in time for the tollow-up meeting at 9:00. Piedmont Fasteners Corporation makes three different clothing fasteners in its manufacturing facility in North Carolina. Data concerning these products appear below VelcroMetalNylon Annual sales volume 106,000 212,000 289,000 S2.00 1.60 S0.80 S1.10 1.00 S0.50 Unit selling price Variable expense per unit Total fixed expenses are $265,000 per year All three products are sold in highly competitive markets, so the company is unable to raise prices without losing an unacceptable numbers of customers. The company has an extremely effcivelean production system, so there are no beginning or ending work in process or finished goods inventories Required 1. What is the companys over-all break-even point in dollar sales? 2. Of the total fixed expenses of $265,000, $23,130 could be avoided if the Velcro product is dropped, $103,200 ir the Metal product is dropped, and S56,100 if the Nylon product is dropped. The remaining fixed expenses of $82,570 consist of common fixed expenses such as administrative salaries and rent on the factory building that could be avoided only by going out of business entirely. a. What is the break-even point in unit sales for each product? b. If the company sells exactly the break-even quantity of each product, what will be the overall profit of the company?
Answer #1

The overall break even sales can be computed by calculating the contribution Margin Ratio Velcro Metal Nylon Total Sales($) 1,65,000 3,00,000 3,40,000 8,05,000 Variable expenses ($) 1,25,000 1,40,000 1,00,000 3,65,000 CM 40,000 1,60,000 2,40,000 4,40,000 Fixed Expenses ($) 4,00,000 Net Operating Income ($) 40,000 CM ratio=Contribution Margin/Sales = 440,000/805,000=0.5466 Dollar sales to Break even=Fixed expenses/CM ratio = 400,000/0.5466=731,800 (approx) 2. The issue is what to do with the common fixed cost when computing the break evens for the individual products. The correct approach is to ignore common fixed costs. If the common fixed costs are used in the computation the break even points would be overstated for individual projects and mangers may drop projects that are profitable. a). The break even points for each product can be computed using the CM approach Velcro Metal Nylon Unit Selling Price ($) 1.65 1.5 0.85 Variable cost per unit 1.25 0.7 0.25 Unit contribution margin(a)($) 0.4 0.8 0.6 Product fixed expenses(b)($) 20,000 80,000 60,000 Unit sales to break even (b/a) 50000 100000 100000 b). Allocation of common fixed expenses on the basis of sales revenue Velcro Metal Nylon Total Sales($) 1,65,000 3,00,000 3,40,000 8,05,000 Percentage of total sales 20.50% 37.27% 42.24% 100% Allocated common fixed expense($) 49,193 89,441 1,01,366 2,40,000 Product fixed expenses 20,000 80,000 60,000 160000 Allocated common and product fixed expenses (a) ($) 69,193 1,69,441